OpenStudy (anonymous):

simplify the complex fraction n-4/n^2-2n-15/n+1/n+3

OpenStudy (anonymous):

@EvanEllis88888

OpenStudy (anonymous):

Simplifying n + -4n2 + -2n + -15n + 1n + 3 = 0 Reorder the terms: 3 + n + -2n + -15n + 1n + -4n2 = 0 Combine like terms: n + -2n = -1n 3 + -1n + -15n + 1n + -4n2 = 0 Combine like terms: -1n + -15n = -16n 3 + -16n + 1n + -4n2 = 0 Combine like terms: -16n + 1n = -15n 3 + -15n + -4n2 = 0

OpenStudy (anonymous):

uhhh what??

OpenStudy (anonymous):

Oh wait i forgot the devison sign :3

OpenStudy (anonymous):

?